OpenStudy (anonymous):

logx - log(x/100) = x

mathslover (mathslover):

Use this to simplify : \(\large{\boxed{\color{blue}{\log(\cfrac{a}{b})=\log a - \log b}}}\)

mathslover (mathslover):

Simplify the second term by using the above property!